12.2 #6

Please lable your answer(s) in bold text.

Evaluate the integral. Check your results by differentiation. (Use C for the constant of integration.) 6x4(3x5 + 7)-3 dx
Answer #1

given integral I = I onu (323+7) de dre here Flul = Gay 3 I = کہر3) - Grey (325 +7) t = 345 +7 Let dt= 1524 de こ ① I dt. z dn